B=1/3 pi x2n solve for N
Answer by stanbon(75887)

B=1/3 pi x2n solve for N
-------
3B = pi*(2n)
-----
2n = (3B)/pi
----
n = (3B)/(2pi)
-----
n = (3/2)(B/pi)
========================
Cheers,
Stan H.
=====================